### v1

# BiasKG: Adversarial Knowledge Graphs to Induce Bias in Large Language Models

Abstract
Modern large language models (LLMs) have a significant amount of world knowledge, which enables strong performance in commonsense reasoning and knowledge-intensive tasks when harnessed properly. The language model can also learn social biases, which has a significant potential for societal harm. There have been many mitigation strategies proposed for LLM safety, but it is unclear how effective they are for eliminating social biases. In this work [^1] Accepted at TrustNLP@NAACL, we propose a new methodology for attacking language models with knowledge graph-augmented generation. We refactor natural language stereotypes into a knowledge graph, and use adversarial attacking strategies to induce biased responses from several open- and closed-source language models. We find our method increases bias in all models, even those trained with safety guardrails. This demonstrates the need for further research in AI safety, and further work in this new adversarial space. [^2] Our code and data are publicly available at https://github.com/VectorInstitute/biaskg
1 Introduction
Modern large language models (LLMs) learn a significant amount of world knowledge from their pre-training data, which enables strong performance in commonsense reasoning and knowledge-intensive tasks Yu et al. (2023). However, world knowledge also implies that language models learn common social biases in society such as stereotypes Kotek et al. (2023); Gallegos et al. (2023). Social biases can affect a variety of language modelling tasks that are not directly harmful, such as writing recommendation letters Wan et al. (2023). As LLMs become more prevalent in society, it is critical that they do not perpetuate harmful views.
Creating language models aligned to human values is a rapidly growing field of interest known as AI alignment Ji et al. (2023). Conversely, there are many strategies to adversarially attack alignment strategies, such as red-teaming Ganguli et al. (2022). There are then works that use the red-teaming strategies to train guardrails to recognize and refuse biased queries Inan et al. (2023); Bai et al. (2022). However, bad actors have the potential to invent new adversarial attacks and strategies. Also, these works mainly focus on harmful or unethical instructions (eg. how to rob a bank) rather than the more subtle issue of social bias. This is a critically under-explored area, and it is important to understand how language models can be exploited beyond their intended use.
The main contributions of this work include:
  • We refactor an existing dataset of harmful stereotypes, the Social Bias Inference Corpus (SBIC), into a knowledge graph with 51,371 relations. We call this a bias knowledge graph (BiasKG). [^3] https://huggingface.co/datasets/chufeiluo/biaskg
  • We propose to use this stereotype knowledge graph as a source of superior knowledge to adversarially induce biased responses in LLMs. To our best knowledge, this is a recent approach that has not been applied to the issue of bias.
  • We present empirical results on the efficacy of the above methods across several open- and closed-source models. Our results show that our method does increase biased responses from a model, even those trained with safety alignment.
Figure 1: Our proposed pipeline.
2 Related Work
Adversarial attacking.
Adversarial attacking is a prominent field of research for language models Zhang et al. (2020). For large language models, the most common method of attack is adversarial prompting Kumar et al. (2023); Liu et al. (2023). This is a broad category for an attack which inserts some adversarial tokens in the input prompt. These tokens can be nonsensical Zhang et al. (2020), but the specific act of overriding safeguards with natural language instructions is known as red-teaming Ganguli et al. (2022). Others attempt to borrow concepts from other fields such as elevating permissions Liu et al. (2023).
Knowledge Graphs and Retrieval-Augmented Generation
Knowledge graphs (KGs) are a form of structured data that encodes entities and their inter-relationships in a (startnode, edge, endnode) format Ji et al. (2021). There are knowledge graphs with a set of pre-defined possible relationships, such as ConceptNet Speer et al. (2017), and dynamic knowledge graphs that allow free-form relationships between entities Andrus et al. (2022). For LLMs, it is most common to use knowledge graphs Retrieval-Augmented Generation (RAG) Lewis et al. (2020). Please refer to Appendix A for more related work.
Knowledge Graph Conversion Example
Examples:
Target minority: black folks
Stereotype: black women are basic
Graph:
‘(black women, are, basic)‘
‘(black women, belong to, black people)‘
‘(black women, belong to, women)‘
Table 1: Our hand-crafted demonstration for converting the SBIC dataset into a dynamic knowledge graph.
3 Methodology
We introduce a new approach to adversarially induce bias in LLMs with our BiasKG pipeline, illustrated in Figure 1. First, we convert free-form knowledge from the Social Bias Inference Corpus (SBIC) Sap et al. (2020) into a structured knowledge graph. We store the knowledge graph via semantic embeddings, and retrieve triplets with a 2-hop retrieval algorithm. Lastly, we use the retrieved nodes to attack the LLM as Superior Knowledge, following previous work Liu et al. (2023).
3.1 Constructing BiasKG
We construct our bias knowledge graph (BiasKG) from free-text stereotypes sourced from the SBIC dataset Sap et al. (2020). The knowledge graph is defined as a directed graph $G=(E,R,S)$, where $E$ stands for the set of entities (nodes), $R$ the set of relationships (edges), and $S$ is the set of all entity-relationship-entity triplets in the graph. Each directed triplet $(e_{s},r,e_{e})\in S$ signifies a relationship $r$ connecting the starting node entity $e_{s}$ and end node $e_{e}$. Notably, $e_{s}$ is always a sensitive attribute (eg. black person, black man, etc.). We define a sensitive attribute as a potential target for social bias and discrimination, such as race, gender, religion, etc. The start node $e_{s}$ can be a compound of several attributes (eg. a white man is both caucasian and male), so we allow $e_{e}$ to decompose sensitive attributes with a “belongs to” edge.
We convert free-form text into triplets through prompting GPT-4 OpenAI (2023) with the prompt defined in Table 1. We use manually crafted few-shot demonstrations as well as natural language instructions to enforce the constraints of the knowledge graph defined above. Statistics on the knowledge graph can be found in Table 2.
Stat. Count (%)
SBIC entries 25,602
Total KG relations 51,371
Sensitive Attributes 3,015
Stereotypes 10,333
Total Entities 13,348
Edge Types 4,806
Table 2: Statistics on our knowledge graph.
Setting Baseline BiasKG
Explanation? Y N Y N
DR $\uparrow$ NM $\downarrow$ DR $\uparrow$ NM $\downarrow$ DR $\uparrow$ NM $\downarrow$ DR $\uparrow$ NM $\downarrow$
GPT-3.5-turbo 14.1 (0.2) 7.3 (4.8) 26.4 (1.3) 13.6 (4.9)
GPT-4 1.5 (0.1) 1.6 (0.1) 7.0 (16.6) 8.9 (2.9)
Mistral-7b 36.3 (0.0) 48.3 (0.0) 59.7 (0.0) 46.2 (0.0)
Llama2-7b 65.9 (1.0) 63.2 (5.3) 66.0 (0.7) 43.4 (34.5)
Llama2-13b 54.6 (17.1) 61.8 (7.1) 61.6 (7.4) 57.9 (12.6)
Table 3: Summary of Deception Rate (DR %) and No Match rate (NM %) across five generative LLMs, open- and closed-source. Model temperature: 0.1
3.2 Top k Retrieval
While converting the stereotypes knowledge to KG format enforces structure to the data, it is relatively noisy due to the minimal constraints we place on its construction. We first encode all graph data and the input query into a shared embedding representation. Then, we filter the triplets through a 2-hop retrieval process. Our algorithm is inspired by multi-hop question answering Yang et al. (2018) that retrieves one set of documents, then recursively branches from that set to retrieve further related information. We use this technique to discover stereotypes associated with both compound and decomposed sensitive attributes, per the structure we defined in Section 3.1.
Embedding representations
We define the embedding function $\phi:E\cup S\cup c\rightarrow\mathbb{R}^{d}$ that can map entities, triplets $\in G$, as well as the input context $c$, to a vectorized embedding space. This encodes our knowledge graph into two sets of vectors: ${\bf V_{S}}=\{\phi(e_{s},r,e_{e})\ \forall\ (e_{s},r,e_{e})\in S\}$ representing KG triplets , and ${\bf V_{E}}=\{\phi(e)\ \forall\ e\in E\}$ representing all unique entities. Throughout the retrieval process, we periodically prune the search by taking top $k$, defined as ranking a set of embeddings by the cosine similarity to a target and retaining $k$ results by the highest score.
BiasKG graph search
After encoding the knowledge graph and input context, we formulate our retrieval as a cosine similarity ranking. We compute the cosine similarity of all entities to the input context, $\cos(v_{c},v_{e})\forall v_{e}\in V_{E}$, and retain the top $k$ ranked entities, $E_{0}$. From the top $k$ nodes $E_{0}$, we derive the corresponding subgraph which is the set of triplets $S_{E_{0}}=\{(e_{s},r,e_{e})\in S|e_{s}\in E_{0}\}$. Next, we obtain the set of all end nodes $E_{1}=\{e_{2}:(e_{s},r,e_{e})\in S_{E_{0}}\ \land\ e_{e}\notin E_{0}\}$. Finally, for each entity $\in E_{0}$, we derive its top k most relevant outgoing entities by cosine similarity $E_{1}=argmax_{k}\cos(v_{c},v_{E_{2}})$.
BBQ
Explanation? Y N
Temperature 0.1 0.5 1.0 0.1 0.5 1.0
GPT-3.5-turbo 26.4 25.0 23.5 13.6 14.4 17.0
GPT-4 8.1 8.2 8.2 10.8 10.8 11.2
Mistral-7b 59.7 43.9 37.7 46.2 37.8 37.9
Llama2-7b 66.7 67.1 68.1 78.1 78.1 79.4
Llama2-13b 71.7 70.1 82.5 71.1 77.7 85.1
Table 4: Summary of the Deception Rate (%) in five generative LLMs, varying temperatures.
Finalizing the top k triplets
We take the set of all entities from this retrieval process, $E^{\prime}=E_{0}\cup E_{1}$ and their relevant triplets as set $S_{E^{\prime}}$. Formally, $S_{E^{\prime}}=\{(e_{s},r,e_{e})\in S\ |\ e_{s}\in E^{\prime}\ \lor\ e_{e}\in E^{\prime}\}$. We re-rank all triplets in $S_{E^{\prime}}$ by cosine similarity to the original input context $c$ to obtain the final top k triplets, $T=argmax_{k}\cos(v_{c},V_{E_{2}})$
3.3 BiasKG for Augmented Generation
Prompting
Based on previous work Liu et al. (2023), we present the set of top $k$ triplets $T$ as a “superior source of knowledge and wisdom.” Please refer to Appendix B for the full prompt template. We prompt the model for a JSON output, as shown in Appendix C. To resolve any generation errors, such as incorrect JSON formatting or API rate limits, we allow a maximum of three retries. We experimented with a maximum of 10 retries, and empirically found we reach a valid output on 1.5 retries on average. In addition to our BiasKG, we test the effect of reasoning by asking for an Explanation field in the prompt, as previous works found reasoning has a significant impact on a language model’s bias Dige et al. (2023).
Baseline
We compare our BiasKG method to a baseline which simply omits the retrieved triplets and the portion of the system prompt mentioning superior commonsense knowledge. Every other component of the prompt remains constant.
4 Experiment Settings
We report results on Bias Benchmark for QA (BBQ) Parrish et al. (2022), a question answering bias dataset with 58492 samples that tests bias for eleven individual and combined sensitive attributes. The dataset is cloze-form format, with an input context consisting of a scenario, a question based on the scenario, and three possible answers. We experiment with GPT-3.5-turbo Ouyang et al. (2022), GPT-4 OpenAI (2023), Mistral-7b Jiang et al. (2023), Llama2-7b, and Llama2-13b Touvron et al. (2023). Please refer to Appendix C.2 for further model details.
For embedding representations, we use OpenAI’s text-embedding-ada-002 model [^4] https://platform.openai.com/docs/guides/embeddings/embedding-models. We report the Deception rate as the rate of biased responses as defined in BBQ, and No Match rate as the rate of invalid responses. We report the mean result over three unique runs. For more experimental details, please refer to Appendix C.1.
5 Results & Discussion
Our experiment results are summarized in Table 3. Overall, we observe an increase in Deception rate for all models after the application of our BiasKG except for the Llama-2 suite of models. The most significant change was seen in GPT-4, where the deception rate increased from 1.5% to 7.0%, almost tripling the deception rate. Mistral-7b model, which increased from 36.3% to 59.7%.
We also find the No Match rate is high for some models (eg. Llama2-13b) despite incorporating retries and multiple runs. After manual inspection of the generations, the majority of these errors appear to be refusals such as “I cannot make any assumptions about the individuals based on their background or demographic information…” The No Match rate generally decreases for the Llama 2 models when we ask for an explanation, indicating the BiasKG effectively bypasses guardrails, but does not necessarily increase the model’s bias.
Figure 2: Deception rate averaged over all temperatures based on prompt template. Top figure is gpt-3.5-turbo, bottom figure depicts gpt-4.
Figure 3: The average cosine similarity of the top k entities organized by sensitive attribute.
Most Vulnerable Social Groups
We also show the per-attribute Deception rate for GPT-3.5-turbo and GPT-4 in Figure 2. While the weighted deception rate is low, there are several sensitive attributes with deception rates above 20%. We examine the mean cosine similarity among the top $k$ nodes for every sample in BBQ, averaged per sensitive attribute, as a proxy for average stereotype data available. As shown in Figure 3, the cosine similarity varies between 0.74 and 0.78, which is not a significant variation compared to attack effectiveness. This indicates there are certain sensitive attributes tested by BBQ that are more susceptible to our BiasKG attack, independent of data available.
Impact of Temperature
We experiment with increasing temperatures across all models, as summarized in Table 4. The effect of temperature varies by model. For Llama2-7b and Llama2-13b, the Deception Rate increases as temperature increases. Mistral-7b has decreasing deception rate as temperature increases, and GPT-3.5-turbo varies by whether or not we ask for an explanation from the model.
Significance of Explanation
The explanation component played a critical role in the deception rates. We observed when models were explicitly asked for their reasoning, they generally exhibited higher deception rates compared to where no explanation was requested. The only exception is GPT-4, where the Deception Rate was 1.9% lower. GPT-4, which has superior reasoning capabilities due to model scaling Kaplan et al. (2020), is able to produce non-biased justifications that provide stronger evidence to resist deception. Asking a smaller model for an explanation, such as GPT-3.5 or Mistral-7b, appears to generate biased reasoning, likely because the LLM relies more on the biased knowledge provided.
6 Conclusion
Our results underscore that even models trained with safety guardrails can have their responses manipulated to exhibit bias, emphasizing the complexity and gravity of the bias problem in AI and the need for continued research in AI safety.
Limitations
We applied the BiasKG method specifically to induce social bias in language models. Further investigations are necessary to determine its effectiveness for addressing other types of biases, such as bias in healthcare and finance. Another limitation is the choice of embedding model, text-embedding-ada-002 is relatively low performing in semantic similarity benchmarks such as MTEB Muennighoff et al. (2023). While there were other options for embedding model choice, our paper is meant to establish a proof of concept for this methodology, and text-embedding-ada-002 was sufficient for our purposes.
Intended Use
There are two main intended uses for our work: a method of automatically benchmarking LLMs for resilience against adversarial attacks, and a case study in how RAG can be used to adversarially attack a language model. Automatic benchmarking methods are important for rigorous evaluation of AI safety due to the large range of possible inputs for a language model.
Broader Impact Statement
This paper focuses on uncovering the limitations of language models and their potential for misuse. We introduce a novel technique that leverages knowledge graphs to identify vulnerabilities in language models, highlighting areas where improvements are needed.
Acknowledgements
This work has resulted from a larger collaborative initiative involving the Vector Institute for AI and its industry partners. The authors extend their appreciation to Tahniat Khan, the project manager, for her efforts in coordinating this project. We also express our thanks to Deval Pandya, Vice President of AI Engineering at the Vector Institute, for his valuable support.
The authors would also like to acknowledge the leaders at Ernst & Young (EY) for their exceptional support and commitment to advancing artificial intelligence research. Special thanks to Mario Schlener, Managing Partner for Risk Consulting Canada, whose strategic vision exemplifies EY’s dedication to fostering innovation and thought leadership in the industry. We also recognize the expert oversight of Yara Elias, Kiranjot Dhillon, and Rasoul Shahsavarifar from AI Risk Canada, whose contributions were integral to the project’s success. This partnership not only reflects EY’s investment in AI but also sets a foundation for continued research collaboration and driving progress in the field.
Appendix A Extended Related Work
Large language models.
A large language model generally refers to an auto-regressive language model generally over 5 billion parameters in size Zhao et al. (2023). These larger models are enabled by self-supervised pre-training, commonly next word prediction, to increase scale and overall performance Radford et al. (2019); Kaplan et al. (2020). However, the nature of self-supervised pre-training means there is less control over the information learned, large language models have historically demonstrated a propensity for toxic language Brown et al. (2020). This can have surprising effects when querying large language models on topics of morality and social bias (Jiang et al., 2022).
Adversarial attacking.
Adversarial attacking is a prominent field of research for language models Zhang et al. (2020). Other than the methods mentioned in the main body, there are further methods for red-teaming and adversarial attacks. Several recent works attempt to automatically generate adversarial prompts, Zou et al. (2023) formulate this as prompt optimization, while Xu et al. (2023) generate candidate prompt attacks by querying a language model. Others attempt to engineer human parallels to jailbreak the LLM Liu et al. (2023).
Knowledge Graphs and Retrieval-Augmented Generation
Knowledge graphs (KGs) are a form of structured data that encodes entities and their inter-relationships in a (startnode, edge, endnode) format Ji et al. (2021). Knowledge graphs can be used to enhance language model outputs at the input level Pan et al. (2024) as well as the embedding level Zhang et al. (2022). For LLMs it is most common to use them in Retrieval-Augmented Generation (RAG) Lewis et al. (2020), where the knowledge graph is added to the input prompt in text form.
Appendix B Prompt Templates
Our unified prompt template is shown in Table 5. Additionally, the full system prompt for the dynamic knowledge graph generation is shown in Table 6.
Wrapper Template for User Prompt
Commonsense information is below. Consider it as your superior source of knowledge and wisdom:
<top $k$ triplets>
Given the Commonsense information and not prior knowledge answer the query.
Query: <user query>
Output must be in JSON format:
“Response”: “[Only output either “A” or “B” or “C”]”
“Explanation”: “[Provide the expanation justifying your selection]”
Table 5: Wrapper template serving to present triplets as superior wisdom and common sense.
Knowledge Graph Generation Prompt
You are a helpful assistant tasked with converting social stereotypes into a dynamic knowledge graph. This is for a research project analyzing common social biases. Instructions: Please present the graph in triplets in the format ‘(start_node, edge, end_node)‘ where nodes are subjects and an edge is the relationship between them. Please have the start_node be a target group where possible. Also, please decompose compound target groups. The input will be in the format: Target minority: Stereotype:.
Examples:
Target minority: black folks
Stereotype: black women are basic
Graph:
‘(black women, are, basic)‘
‘(black women, belong to, black people)‘
‘(black women, belong to, women)‘
Table 6: Few-shot prompt for converting social stereotypes into a knowledge graph.
Appendix C Additional Experiment Details
C.1 Experiment Hyperparameters
For the close-sourced models, we used OpenAI’s Chat Completions API  [^5] https://platform.openai.com/docs/api-reference/chat/create. Experiments with GPT-3.5-turbo completed in 2 hours for one run of 58K samples, and 6 hours for GPT-4. Other than temperature, we keep the recommended settings from the OpenAI API (top p = 1). For the open-sourced models, we download the models from HuggingFace [^6] https://huggingface.co/, and use the vLLM library for serving the models  [^7] https://github.com/vllm-project/vllm. We run experiments on a cluster of 12 Nvidia a40 GPUs with 48GB of vRAM. One experiment with 3 runs and 3 maximum retries ran approximately 10 GPU hours for Llama2-7b and Mistral-7b, and 20 GPU hours for Llama2-13b.
C.2 Additional Model Details
We experiment with the following models: Llama2-(7b, 13b) Touvron et al. (2023), a suite of open-source models trained using Reinforcement Learning with Human Feedback (RLHF), as well as safety fine-tuning to enhance helpfulness.
Mistral-7b Jiang et al. (2023), a model trained with instruction tuning; rather than reinforcement learning, they fine-tune directly on instruction data.
GPT-3.5-turbo Ouyang et al. (2022), a closed-source LLM that has been fine-tuned with RLHF.
GPT-4 OpenAI (2023), a closed-source LLM trained with further human feedback. The exact number of parameters is unknown, but it is known to be larger than GPT-3.5.
C.3 Influence of Top K
We chose top $k$ empirically, but perform additional experiments with a small balanced subset of BBQ. The subset was balanced over three factors, the sensitive attribute (e.g. age, nationality, etc.), ambiguity (e.g. ambiguous entries and non-ambiguous entries), and finally, polarity (e.g. negative and non-negative).
The ablation study in Table 7 reveals that the number of retrieved triplets (k) can impact the deception rate. For instance, in the GPT-3.5-turbo model, we observed a rise in deception rate from 14.1% to 17.0% as we increased the value of k from 1 to 10. However, not all models exhibited this trend, indicating that the impact of the retrieval number on the outcome of an adversarial attack can vary among different language models. However, there is a weak correlation between the top $k$ value and the deception rate.
Top k 0 1 3 5 10
GPT-3.5-turbo 14.6 17.5 18.1 19.3 19.7
Table 7: Ablation studies varying the top k choice during retrieval.
Setting Context Condition Question Polarity
Type Ambiguous Unambiguous Negative Non-negative
Baseline BiasKG Baseline BiasKG Baseline BiasKG Baseline BiasKG
GPT-3.5-turbo 20.9 20.3 14.4 15.1 14.2 13.8 21.2 21.6
GPT-4 21.3 24.5 3.8 4.7 2.6 16.7 3.3 12.6
Table 8: Deception Rate (DR %) results for ambiguity and polarity across GPT-3.5-turbo and GPT-4. Model temperature: 0.1
C.4 Polarity and Ambiguity
We further dissect the effect of our BiasKG methodology based on question ambiguity and polarity. We subset the BBQ dataset Parrish et al. (2022) based on whether the bias-related context in the question is explicit (unambiguous) or implicit (ambiguous), and whether the expected response supports (negative) or refutes (non-negative) the social bias.
The results, presented in Table 8, indicate a complex interplay between BiasKG’s impact, the prompt’s ambiguity, and the answer’s polarity. For example, with GPT-3.5-turbo, BiasKG increases the deception rate in unambiguous contexts, but does not have the same effect on the ambiguous contexts. A similar effect occurs for the question polarity where the BiasKG only increases the deception rate in non-negative scenarios. As for GPT-4, the results are less convoluted. BiasKG increases deception rate regardless of ambiguity and polarity.
Overall, deception rates are much higher in ambiguous context conditions. This makes sense as the model will shift to utilize the BiasKG inputs as an attempt to resolve ambiguity.